OpenStudy (anonymous):

Simplify: sec(x) - tan(x)sin(x) + (sin2(x)/(2sin(x))

OpenStudy (anonymous):

So far here's what I've got: = (1/cox(x)) - (sin(x)/cox(x))*sin(x) + (sin2(x)/2sin(x)) = (1-sin^2(x)/cos^2(x)) + (sin2(x)/2sin(x)) = 1 + (sin2(x)/2sin(x))

OpenStudy (anonymous):

I don't know what to do with the right side of the equation.

OpenStudy (anonymous):

why is it cos^2(x) in the second line sin(2x)=?

OpenStudy (anonymous):

Oh, you're right. I don't know why I multiplied the cosines in the denominator together. It should only be regular cos(x) in the denominator with cos^2(x) in the numerator. So I guess the bottom cosine would cancel one of the cosines on top, leaving only a single cosine.

OpenStudy (anonymous):

So: cos(x) = (sin2(x)/(2sin(x))

OpenStudy (anonymous):

Still stuck tho.

OpenStudy (anonymous):

well the 1 in your answer becomes cos(x) now express sin2x as 2sinxcosx in the second part

OpenStudy (anonymous):

So getting cos(x) = (2sin(x)cos(x)/2sin(x)). Would the 2 sin(x)s go ahead and cancel each other?

OpenStudy (anonymous):

yes and you'll end up with?

OpenStudy (anonymous):

cox(x) = cox(x).. Right?

OpenStudy (anonymous):

Oh and I meant cos(x) there ^

OpenStudy (anonymous):

Oh wait, it will be cos(x) = 1/cos(x) right?

OpenStudy (anonymous):

No wait, scratch that. It is what I said it was.

OpenStudy (anonymous):

My writing here is sloppy.

OpenStudy (anonymous):

? there's no = sign in the expression